# TokaMark: A Comprehensive Benchmark for MAST Tokamak Plasma Models

### v1

@@ 2.2. Tasks Structural Taxonomy
Each task consists of one or more input signals and one or more output signals also referred to as targets. Inputs typically include both sensor signals (e.g., magnetics, radiatives, kinetics) and actuator signals (e.g., voltages, fueling), while outputs include either sensor signals or derived quantities like equilibrium reconstructions.
@@
  • Reconstruction. Given a set of sensors signals $A$ over the interval $t_{0}-\Delta_{input},t_{0}]$, the goal is to reconstruct a related set of signals $B$ over the same interval.
  • Autoregressive (AR) Forecasting. Given sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ together with actuator trajectories over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, predict future values of the same signal set over $t_{0},t_{0}+\Delta_{output}]$.
  • Reconstructive (RC) Forecasting. Using sensors signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ and actuators over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, forecast a set of related outputs $B$ over $t_{0},t_{0}+\Delta_{output}]$ where $B$ may contain signals as in $A$.
We further distinguish tasks by their temporal dependency structure: Markovian tasks require only a short input window to make forecasts, reflecting fast dynamics, whereas Non-Markovian (NM) tasks require substantially longer input histories. The latter arise in phenomena governed by slower transport processes or evolving stability conditions, where even short-term predictions depend on extended plasma context.
Finally, TokaMark emphasizes self-supervised and weakly supervised task formulations whenever possible, reducing reliance on expert-labeled targets. This design was taken to encourage the development of AI models capable of discovering transferable plasma representations.
@@ 2.3. Data-driven Challenges
To complete the description of the FAIR-MAST dataset we highlight three core data-centric challenges that AI systems must overcome in order to resolve tasks in TokaMark. These challenges arise directly from the heterogeneous, multi-instrument nature of tokamak diagnostics and the realities of operating large fusion experiments.
  • Multi-fidelity. Diagnostic systems operate at widely varying sampling rates, ranging from a few hundred hertz to several hundred kilohertz. Resampling to a common high-frequency time base is computationally expensive and often unnecessary, while down-sampling to the lowest frequency discards information critical for resolving fast plasma phenomena. Effective models must therefore integrate and represent multi-rate temporal data without losing fidelity.
  • Multi-modality. Signals differ not only in physical meaning, units, numerical ranges, but also in structural forms of the corresponding data tensors. The dimensionality of the latter can be from 1 to 3 with the number of channel in each non-temporal dimension vary from 1 up to 170. AI systems must be able to fuse these diverse modalities into coherent representations suitable for downstream tasks.
  • Missing data. As with any experimental dataset, missing information is common. Entire signals may be absent for a given shot due to hardware issues. Individual signals may also contain missing time segments, for example due to limited acquisition windows or diagnostic failures. Naïvely discarding shots or windows with missing components wastes valuable examples and can introduce distributional bias. Robust approaches must therefore handle incomplete signals and irregular temporal coverage in a principled manner.
3. MAST Benchmark
In this section, we provide a description of the main components of TokaMark: tasks, data preparation and evaluation protocol.
@@ 3.1. Downstream Tasks
Task Input Sensors Input Actuators Outputs Type Input Window Output Window
Group 1 1-1 Magnetics, Currents Shape parameters, $\beta$ metrics Reconstruction 5 ms 5 ms
@@ 1-3 Same as 1-1 Flux map Reconstruction 5 ms 5 ms
Group 2 2-1 Magnetics, Currents Voltages, NBI Power Currents, $\beta$ metrics, Shape parameters RC Forecasting 5 ms 25 ms
@@ 3-2 Thomson scattering, Radiatives Same as 3-1 Radiatives AR Forecasting 5 ms 50 ms
3-3 Magnetics, Currents, Interferometer Same as 3-1 Thomson scattering, $\beta$ metrics RC Forecasting, NM history 5 ms
@@ 4-5 Same as 4-1 Same as 4-1 Mirnov diagnostics AR Forecasting, NM history 100 ms
For the TokaMark we constructed a suite of 14 tasks designed to evaluate the core capabilities expected of AI models for real-world fusion experiments: learning unified representations from heterogeneous diagnostics, temporal reasoning across both fast and slow dynamics, robustness to incomplete state information, and generalisation across operating regimes. Those tasks span across diverse timescales, modalities, and predictive objectives.
We arranged the tasks into 4 broad groups that represent some of the major modeling challenges an AI system must master: instantaneous reconstruction, short‑term magnetics dynamics, slow transport‑driven profile evolution, and long‑range forecasting of rare but safety‑critical events and MHD activity. Table 2 summarizes definitions of tasks and groups.
@@ 3.1.1. Group 1: Equilibrium Reconstruction
The tasks in group 1 define a suite of reconstruction problems where the objective is to infer plasma equilibrium – its shape, boundaries, and various properties – from instantaneous magnetic measurements. The targets of the three tasks in this group span multiple modalities from scalar parameters (task_1-1) to full plasma contour (task_1-2) and two-dimensional representations of the poloidal magnetic flux (task_1-3). The inputs of the tasks are identical and use standard magnetic diagnostics and actuator currents.
Equilibrium reconstruction is a foundational operation in tokamak experiments, performed routinely after every plasma discharge and forming the basis of essentially all downstream analysis. Traditionally, equilibrium is reconstructed by solving an inverse boundary-value problem governed by the Grad–Shafranov equation that describes the balance of magnetohydrodynamic force in axisymmetric tokamak plasmas. These solvers require multiple iterations for convergence and rely on parameterized assumptions about plasma profiles, making it challenging to deploy them in real time.
Group 1 tasks evaluate whether AI models can infer equilibrium structure directly from raw diagnostics, providing fast, numerics-free surrogates for a core analysis task that could underpin both offline interpretation and real-time control applications.
@@ 3.1.2. Group 2: Magnetics Dynamics
The tasks in Group 2 address short-timescale forecasting of magnetic signals, coil currents, and equilibrium evolution in response to applied actuator commands. Conceptually, these tasks parallel those in Group 1, but shift from static reconstruction to sequence-to-sequence forecasting with actuators signals now playing an essential role in the tasks inputs. Complexity increases in the three tasks in the group: from forecasting scalar quantities - such as plasma current evolution - to joint prediction of 2D equilibrium geometry over short horizons.
At these timescales, plasma evolution is dominated by inductive coupling between active coils, passive conducting structures, and the plasma current, together with the local response of the equilibrium to perturbations. These dynamics govern how the plasma reacts to control actions during a discharge and are central to plasma position, shape, and current control. While these effects are routinely modeled offline, a model capable of achieving real time latency would enable improved control policies.
Group 2 probes whether models can learn the effective plasma– machine response encoded in coupled circuit dynamics and plasma response. This would directly target capabilities required for closed-loop control, scenario planning, and digital twin applications.
@@ 3.1.3. Group 3: Profile Dynamics
Group 3 tasks focus on modeling the temporal evolution of kinetic profiles - primarily electron density and temperature - and on forecasting diagnostic signals associated with confinement-mode transitions. The three tasks in this group include short-horizon forecasting (task_3-1 and task_3-2) and pseudo-real-time reconstruction using partial, multi-rate diagnostic inputs (task_3-3).
Profile evolution is governed by transport physics acting on particle and energy balance, introducing slower characteristic timescales and intrinsic memory effects compared to magnetic dynamics. These processes control energy confinement and overall plasma performance and are central to understanding transport, optimising heating schemes, and predicting access to high-performance regimes. In practice, profile measurements are often sparse, delayed, or unavailable in real time.
The main goal of Group 3 tasks is to assess AI models ability to integrate incomplete diagnostic information over time to infer latent plasma state variables. This would support applications ranging from post-shot analysis, to integrated scenario design and real-time performance monitoring and control.
@@ 3.1.4. Group 4: MHD activity
Finally, the tasks in Group 4 comprise long-horizon forecasting of thermal quenches (task_4-1 and task_4-2), vertical displacement events (task_4-3), current quenches (task_4-4) and MHD activity leading to Locked Modes (task_4-5). These tasks are focused on early detection of plasma instability and disruptions precursors that emerge across multiple diagnostic modalities. These tasks require processing of rich, high frequency inputs–including magnetics, radiative diagnostics, and time–frequency MHD signatures—over extended temporal windows, exceeding well beyond immediate system response times.
The onset of MHD instabilities and disruptions is tightly linked to the evolving equilibrium and current distribution, even when the nonlinear dynamics of the instability itself are difficult to model explicitly. Disruption avoidance and mitigation are critical operational requirements for present-day tokamaks and future reactors, directly affecting machine protection. In practice, early-warning systems rely on detecting subtle precursors distributed across multiple diagnostics and evolving over extended time windows.
Group 4 tasks probe models abilities to integrate long-range temporal context and multi-modal information to anticipate loss-of-control events, testing capabilities that are essential for safe and reliable fusion operation.
@@ 3.2. Data Preparation
From the FAIR-MAST dataset we take all 11,573 available shots and extract the data for 39 signals required by out task definitions. Before being used for model training and evaluation, these signals undergo through a standardized preprocessing pipeline to ensure consistent formatting and alignment.
Data split. The benchmark dataset is divided into disjoint training, validation, and test subsets. To avoid information leakage across sets, the split is performed at the shot level representing independent experiments and employs random sampling technique. The split ratio is 80% / 10% / 10%, and all hyperparameter tuning is conducted exclusively on the training and validation sets. The test set is held out and used only for the final performance evaluation.
Window segmentation. Shot-level signals data is segmented into multiple input and output windows. The lengths of those windows are task dependent and specified in Table 2. Windows are extracted using a sliding-window approach with a stride of $0.001$ seconds between consecutive windows. Each input window is paired with its corresponding output window and treated as an independent window-level sample. We note that signals are used at their original sampling rates, and no resampling or imputation is performed.
Handling missing information. We impose additional constraints on the test set to ensure meaningful evaluation. Windows in which all input signals are entirely absent or composed solely of NaN values are excluded. Moreover, all output signals must be fully available over the entire prediction horizon.
@@ 3.3. Benchmark Evaluation
As part of the TokaMark, we introduce a hierarchical evaluation protocol for assessing model performance in a manner that aligns with scientific objectives and the structure of FAIR-MAST data. The design explicitly separates three levels of the hierarchy: (i) signals (individual physical quantities), and (ii) tasks (well-defined scientific goals) , and (iii) groups (broader physical objectives). This ensures that the benchmark assess both low-level prediction quality and high-level scientific utility. For low-level evaluation, the hierarchy provides granular signal-level insights that help diagnose which physics regimes the model captures or fails to capture. For models trained jointly across multiple tasks, the same hierarchy provides modular evaluations that highlight generality and transfer properties across tasks.
@@
Windows are containers of the equal per signal size storing $N_{k}$ samples. We compute a window-level Root-Mean-Square error error as:
@@
Signals serve as first level of the evaluation hierarchy. For each signal $k$ containing $M_{k}$ windows, we aggregate all window errors via the signal-level Root-Mean-Square error and normalize it with the empirical standard deviation $\sigma_{k}$ computed for signal $k$ across all evaluation shots:
(2) $\mathrm{RMSE}_{k}=\sqrt{\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\mathrm{RMSE}_{k,i}^{2}},\quad\mathrm{NRMSE}_{k}=\frac{\mathrm{RMSE}_{k}}{\sigma_{k}}$
This yields a dimensionless quantity that expresses prediction error relative to the natural variability of the target making the metric comparable across signals. An error of $\mathrm{NRMSE}_{k}=1$ corresponds to a model no better than approximating the signal by its mean, while $\mathrm{NRMSE}_{k}<1$ indicates predictive value. We also note that $\mathrm{RMSE}_{k}$ and $\mathrm{NRMSE}_{k}$ are computed for a given task and shot.
Tasks. Task-level errors $\tilde{e}_{t}$ combine $K_{t}$ normalized output signal errors into a single score using a uniform mean, allowing one to quantify model performance on the scientific objective as a whole:
@@
Shots are the final step of the aggregation sequence. They represent distinct plasma discharges and therefore provide natural boundaries for independent experimental realizations. We report per-signal and per-task shot-level averaged scores correspondingly defined as
(4) $\mathrm{SignalScore}=\frac{1}{S}\sum_{s=1}^{S}\mathrm{NRMSE}_{s,k},\quad\mathrm{TaskScore}=\frac{1}{S}\sum_{s=1}^{S}\tilde{e}_{s,t}$
Here, $S$ is the total number of shots in the test set used for evaluation. Finally, the group level score is taken as an average of the corresponding task scores.

### v2

@@ 2.2. Structural Taxonomy of Tasks
Each task consists of one or more input signals, and one or more output signals also referred to as targets. Inputs typically include both diagnostic signals (e.g., magnetics, radiatives, kinetics) and actuator signals (e.g., voltages, fueling), while outputs include either diagnostic signals or derived quantities like equilibrium reconstructions.
@@
  • Reconstruction. Given a set of diagnostic signals $A$ over the interval $t_{0}-\Delta_{input},t_{0}]$, the goal is to reconstruct a related set of signals $B$ over the same interval.
  • Autoregressive (AR) Forecasting. Given diagnostic signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ together with actuator trajectories over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, predict future values of the same diagnostic signals $A$ over $t_{0},t_{0}+\Delta_{output}]$.
  • Reconstructive (RC) Forecasting. Using diagnostic signals $A$ over $t_{0}-\Delta_{input},t_{0}]$ and actuators over $t_{0}-\Delta_{input},t_{0}+\Delta_{output}]$, forecast a set of related outputs $B$ over $t_{0},t_{0}+\Delta_{output}]$ where $B$ may contain signals as in $A$.
We further distinguish tasks by their temporal dependency structure: Markovian tasks require only a short input window to make forecasts, reflecting fast dynamics, whereas Non-Markovian (NM) tasks require substantially longer input histories even for short-term predictions.
@@ 2.3. Data-driven Challenges
To complete the description of FAIR-MAST data, we highlight three core data-centric challenges that AI systems must overcome in order to resolve tasks in TokaMark. These challenges arise directly from the heterogeneous, multi-instrument nature of tokamak diagnostics and the realities of operating large fusion experiments.
Multi-fidelity. Diagnostic systems operate at varying sampling rates, ranging from a few hundred hertz to several hundred kilohertz. Resampling to a common high-frequency time base is computationally expensive and often unnecessary, while down-sampling to the lowest frequency discards information critical for resolving fast plasma phenomena. Effective models must therefore integrate and represent multi-rate temporal data without losing fidelity.
Multi-modality. Signals differ not only in physical meaning, units, and numerical ranges, but also in structural forms of the corresponding data tensors. Their dimensionality can be from 1 to 3, with the number of channel in each non-temporal dimension varying from 1 up to 170. AI systems must be able to fuse these modalities into coherent representations suitable for downstream tasks.
Missing data. As with any experimental dataset, missing information is common. Entire signals may be absent for a given shot due to hardware issues. Individual signals may also contain missing time segments, for example due to limited acquisition windows or diagnostic failures. Naively discarding shots or windows with missing components wastes valuable examples and can introduce distributional bias. Robust approaches must therefore handle incomplete signals and irregular temporal coverage.
3. TokaMark: A MAST Benchmark
In this section, we provide a description of the main components of TokaMark: tasks, data preparation, and evaluation protocol.
@@ 3.1. Downstream Tasks
Task Input Diagnostics Input Actuators Outputs Type Input Window Output Window
Group 1 1-1 Magnetics, Currents Shape parameters, $J_{\text{tor}}$ metrics Reconstruction 5 ms 5 ms
@@ 1-3 Same as 1-1 Flux map Reconstruction 5 ms 5 ms
Group 2 2-1 Magnetics, Currents Voltages, NBI Power Currents, $J_{\text{tor}}$ metrics, Shape parameters RC Forecasting 5 ms 25 ms
@@ 3-2 Thomson scattering, Radiatives Same as 3-1 Radiatives AR Forecasting 5 ms 50 ms
3-3 Magnetics, Currents, Interferometer Same as 3-1 Thomson scattering, $J_{\text{tor}}$ metrics RC Forecasting, NM history 5 ms
@@ 4-5 Same as 4-1 Same as 4-1 Mirnov diagnostics AR Forecasting, NM history 100 ms
In TokaMark, we assembled a set of 14 tasks spanning across diverse timescales, modalities, and predictive objectives. We arranged them into 4 broad groups that represent some of the major modeling challenges arising in real-world fusion experiments. The 4 groups include: instantaneous reconstruction, short‑term magnetics dynamics, slow transport‑driven profile evolution, and long‑range forecasting of rare but safety‑critical events and MHD activity. Table 2 summarizes definitions of tasks and groups.
@@ 3.1.1. Group 1: Equilibrium Reconstruction
The tasks in Group 1 define a suite of reconstruction problems where the objective is to infer plasma equilibrium—its shape, boundaries, and various properties—from instantaneous magnetic measurements. The targets of the three tasks in this group span multiple modalities from scalar parameters (Task 1-1) to full plasma contour (Task 1-2) and two-dimensional representations of the poloidal magnetic flux (Task 1-3). The inputs of the tasks are identical and use magnetic diagnostics and actuator coil currents.
Equilibrium reconstruction is a foundational operation in tokamak experiments, performed routinely after every plasma discharge and forming the basis of essentially all downstream analysis. Traditionally, equilibrium is reconstructed by solving an inverse boundary-value problem governed by the Grad–Shafranov equation that describes the balance of magnetohydrodynamic force in axisymmetric plasmas. These solvers, like EFIT++ (Appel and Lupelli, 2018), require multiple iterations for convergence and rely on parameterized assumptions about plasma profiles, making it challenging to deploy them in real time.
Group 1 tasks evaluate whether AI models can infer equilibrium structure directly from raw diagnostics, providing fast, numerics-free surrogates that could underpin both offline interpretation and real-time control applications.
@@ 3.1.2. Group 2: Magnetics Dynamics
The tasks in Group 2 address short-timescale forecasting of magnetic signals, coil currents, and equilibrium evolution in response to applied actuator commands. Conceptually, these tasks parallel those in Group 1, but shift from static reconstruction to sequence-to-sequence forecasting with actuators signals now playing an essential role in the tasks inputs. Complexity increases in the three tasks in the group: from forecasting scalar quantities—such as plasma current evolution—to joint prediction of 2D equilibrium geometry over short horizons.
At these timescales, plasma evolution is dominated by inductive coupling between active coils, passive conducting structures, and the plasma current, together with the local response of the equilibrium to perturbations. These dynamics govern how the plasma reacts to magnetic control actions during a discharge and are central to plasma position, shape, and current control. While these effects are routinely modeled offline, a model capable of achieving real-time latency would enable improved control policies.
Group 2 probes whether models can learn an effective description of the plasma response to the coupled magnetic dynamics of the active coils and vessel. This would directly target capabilities required for closed-loop control, scenario planning, and digital twin applications.
@@ 3.1.3. Group 3: Profile Dynamics
Group 3 tasks focus on modeling the temporal evolution of kinetic profiles—primarily, electron density and temperature—and on forecasting diagnostic signals associated with confinement-mode transitions. The three tasks in this group include short-horizon forecasting (Task 3-1 and Task 3-2) and pseudo-real-time reconstruction using partial, multi-rate diagnostic inputs (Task 3-3).
Profile evolution is governed by transport physics acting on particle and energy balance, introducing slower characteristic timescales and intrinsic memory effects compared to magnetic dynamics. These processes control energy confinement and overall plasma performance and are central to understanding transport, optimizing heating schemes, and predicting access to high-performance regimes. In practice, profile measurements are often sparse, delayed, or unavailable in real time.
The main goal of Group 3 tasks is to assess AI models’ ability to integrate incomplete diagnostic information over time to infer latent plasma state variables. This would support applications ranging from post-shot analysis, to integrated scenario design and real-time performance monitoring and control.
@@ 3.1.4. Group 4: MHD activity
Finally, the tasks in Group 4 comprise long-horizon forecasting of thermal quenches (Task 4-1 and Task 4-2), vertical displacement events (Task 4-3), current quenches (Task 4-4), and MHD activity leading to Locked Modes (Task 4-5). These tasks are focused on early detection of plasma instability and disruptions precursors that emerge across multiple diagnostic modalities. These tasks require processing of rich, high-frequency inputs—including magnetics, radiative diagnostics, and time–frequency MHD signatures—over extended temporal windows.
The onset of MHD instabilities and disruptions is tightly linked to the evolving equilibrium and current distribution, even when the nonlinear dynamics of the instability itself are difficult to model explicitly. Disruption avoidance and mitigation are critical operational requirements for present-day tokamaks and future reactors, directly affecting machine lifespan. In practice, early-warning systems rely on detecting subtle precursors distributed across multiple diagnostics and evolving over extended time windows.
Successful performance in Group 4 tasks would demonstrate the ability to integrate long-range temporal context and multi-modal information, allowing the models to anticipate loss-of-control events. This capabilities are essential for safe and reliable fusion operation.
3.1.5. Input and output window length
For fast magnetic and actuator dynamics (Group 2, tasks 3-1 and 3-2), the observed diagnostics provide a sufficient description of the system state, such that short input windows are often adequate to predict near-term evolution. In contrast, profile evolution (task 3-3), confinement transitions, and MHD activity (Group 4) depend on latent plasma state variables—such as current and pressure profiles—that are only partially and indirectly observed. For these processes, accurate prediction requires integrating information over extended time intervals, making long input windows essential. By explicitly varying the temporal context required across tasks, the benchmark probes a model’s ability to learn both short-timescale system dynamics and long-range temporal dependencies arising from unobserved physics.
@@ 3.2. Data Preparation
We take all 11,573 available shots from FAIR-MAST data and extract the data for 39 signals required by our task definitions. Before being used for model training and evaluation, these signals undergo a standardized preprocessing pipeline to ensure consistent formatting and alignment.
Data split. The benchmark dataset is divided into disjoint training, validation, and test subsets. To avoid information leakage across sets, the split is performed at the shot level representing independent experiments, and employs random sampling technique. The split ratio used is 80%/10%/10%, and all hyperparameter tuning is conducted exclusively on the training and validation sets. The test set is held out and used only for the final performance evaluation.
Window segmentation. Shot-level signals data is segmented into input and output windows. The lengths of those windows are task-dependent as specified in Table 2. Windows are extracted using a sliding-window approach with a stride of $0.001$ seconds between consecutive windows. Each input window is paired with its corresponding output window and treated as an independent window-level sample. We note that signals are used at their original sampling rates, and no resampling or imputation is performed.
Handling of missing information. We impose additional constraints on the test set to ensure meaningful evaluation. Windows in which all input signals are entirely absent or composed solely of NaN values are excluded. Moreover, all output signals must be fully available over the entire prediction horizon.
@@ 3.3. Benchmark Evaluation
As part of TokaMark, we introduce a hierarchical evaluation protocol for assessing model performance which explicitly separates three levels of the hierarchy: (i) signals (individual physical quantities), (ii) tasks (well-defined scientific goals), and (iii) groups (broader physical objectives). This provides both, signal-level insights that help diagnose which physics regimes the model captures and higher-level scientific utility assessment.
@@
Windows are containers of the equal per-signal size storing $N_{k}$ samples. We compute a window-level Root-Mean-Square error error as:
@@
For each signal $k$ for a given task and shot and containing $M_{k}$ windows, we aggregate all window errors via the single-shot signal Root-Mean-Square error and normalize it with the empirical standard deviation $\sigma_{k}$ computed for signal $k$ across all evaluation shots:
(2) $\mathrm{RMSE}_{k}=\sqrt{\frac{1}{M_{k}}\sum_{i=1}^{M_{k}}\mathrm{RMSE}_{k,i}^{2}},\quad\mathrm{NRMSE}_{k}=\frac{\mathrm{RMSE}_{k}}{\sigma_{k}},$
which yields a dimensionless quantity that expresses prediction error relative to the natural variability of the target, making the metric comparable across signals. An error of $\mathrm{NRMSE}_{k}=1$ corresponds to a model no better than approximating the signal by its mean, while $\mathrm{NRMSE}_{k}<1$ indicates predictive value.
Concerning tasks, task-level errors $\tilde{e}_{t}$ combine $K_{t}$ normalized output signal errors into a single score using a uniform mean, allowing one to quantify model performance on the scientific objective as a whole:
@@
Shots represent independent experimental realizations, and therefore provide the final step of aggregation. We report signal-level and task-level errors as NRMSE based aggregated across shots quantities correspondingly defined as:
(4) $\text{Signal}_{\mathrm{NRMSE}}=\frac{1}{S}\sum_{s=1}^{S}\mathrm{NRMSE}_{k(s)},\quad\text{Task}_{\mathrm{NRMSE}}=\frac{1}{S}\sum_{s=1}^{S}\tilde{e}_{t(s)}.$
Here, $S$ is the total number of shots in the test set used for evaluation. Finally, the $\text{Group}_{\mathrm{NRMSE}}$ score is taken as an average of the corresponding $\text{Task}_{\mathrm{NRMSE}}$ scores.